# The Bayesian Geometry of Transformer Attention

### v2

@@ 7.3. Mechanistic Interpretability and Attention Geometry
Mechanistic studies of transformers have revealed specialized attention heads for induction, copying, and retrieval (Elhage et al., 2021; Nanda et al., 2023). Other work has examined QKV spaces, circuit decomposition, and sparse structures that arise during training (Olsson et al., 2022). These studies provide qualitative and circuit-level insight into model behaviors.
Our contribution is to link these geometric structures directly to Bayesian inference in a setting where the posterior is known. We show that keys form near-orthogonal hypothesis axes, queries sharpen onto feasible hypotheses across depth, and value representations unfurl into a one-dimensional entropy manifold. This connects mechanistic interpretability to probabilistic computation in a rigorous way: the internal geometry needed for Bayesian reasoning becomes directly visible.
7.4. Architectural Comparisons
Alternative sequence models—-state-space architectures (Gu et al., 2022; Gu and Dao, 2023), convolutional variants (Poli et al., 2023), and deep MLPs—- often match transformers in perplexity on natural text. But perplexity conflates modeling and inference capability. Our results provide a finer test: whether an architecture can reproduce an analytic Bayesian posterior under strict non-memorization constraints. The capacity-matched MLP controls clarify that attention-based routing is essential for Bayesian structure learning in context, at least relative to flat feed-forward architectures.
@@ 7.5. Training Dynamics
Finally, concurrent work analyzes the gradient dynamics that create these structures during training (Aggarwal et al., 2025). They show that attention and value updates follow coupled laws that produce a stable routing frame and a progressively refined value manifold. Our empirical findings align with this picture: attention stabilizes early, while value vectors continue to encode the posterior with increasing resolution. Together, these perspectives connect the optimization trajectory to the geometric structure that implements Bayesian inference.
@@ 8. Limitations and Future Work
Bijections and HMMs capture essential elements of Bayesian computation—-discrete elimination and recursive state tracking—-but they represent only a narrow slice of the inference problems encountered by large language models. Future wind tunnels could incorporate richer latent-variable structures, including Kalman filtering, hierarchical Bayesian models, or causal graphical models, all of which have closed-form posteriors and allow precise verification.
@@
Although the hypothesis spaces in both tasks are large enough to prevent memorization, their representational dimensionality is modest (e.g., five hidden states in HMMs). Larger systems with high-dimensional latent variables would test whether the geometric mechanisms we observe—-orthogonal hypothesis axes, progressive Q–K sharpening, and value-manifold refinement—-scale smoothly with dimensionality.
@@
The experiments here use standard transformers. It remains unclear whether alternative architectures—-state-space models, deep MLPs with more sophisticated gating, or hybrid recurrent-attention systems—-can form comparable Bayesian manifolds. Wind-tunnel evaluations could provide a principled benchmark for comparing architectures in terms of inference fidelity rather than perplexity alone.
@@
A notable empirical phenomenon is the frame—-precision dissociation: attention maps stabilize early while value manifolds continue to unfurl and refine posterior precision. A systematic study of these phases—-how early the frame forms, how quickly precision improves, and how these dynamics depend on depth, width, and data complexity—-could lead to a more general theory of representation formation in transformers.
@@ 9. Conclusion
We introduced Bayesian wind tunnels—-controlled experimental settings with analytic posteriors and combinatorially large hypothesis spaces—-to test whether transformers genuinely implement Bayesian inference rather than merely mimicking it. Across two fundamentally different inference problems, discrete bijection elimination and sequential state tracking in Hidden Markov Models, small transformers converge to the exact Bayesian posterior with sub-bit calibration error, even at sequence lengths well beyond those seen in training. Capacity-matched MLPs fail catastrophically in both settings, demonstrating that this behaviour arises from the geometry of attention rather than model size or optimization.
Geometric diagnostics provide a unified explanation. Keys form an approximately orthogonal basis over hypotheses; queries progressively align with the feasible region of that basis; and value vectors organize along a low-dimensional manifold parameterized by posterior entropy. Training sculpts this manifold: attention patterns stabilize early, while value representations continue refining posterior precision—-a frame—-precision dissociation predicted by concurrent gradient-dynamics analysis. These mechanisms together implement the essential components of Bayesian conditioning: binding, elimination, and refinement, expressed as a sequence of structured linear transformations across depth.
The wind-tunnel regime is intentionally simplified, but it establishes a clear lower bound: if a model cannot implement Bayes in settings where the posterior is known and memorization is impossible, it cannot do so in natural language. Conversely, our results show that transformer geometry is sufficient for exact Bayesian inference when the task permits verification. This provides a principled foundation for studying approximate reasoning in larger models and offers concrete, testable predictions—-orthogonal hypothesis axes, progressive Q–K sharpening, and value-manifold structure—-for analysing pretrained LLMs.
Transformers succeed here because their architecture furnishes the right inductive bias—not because of scale: residual streams that carry evolving belief states, attention that routes information selectively, and feed-forward layers that implement local Bayesian updates. Together, these components carve a Bayesian manifold inside the model’s representation space. Understanding how this manifold emerges, scales, and ultimately degrades in real-world language remains an important direction for future work.

### v3

@@ 7.3. Mechanistic Interpretability and Attention Geometry
Mechanistic studies of transformers have revealed specialized attention heads for induction, copying, and retrieval (Elhage et al., 2021; Nanda et al., 2023). Other work has examined QKV spaces, circuit decomposition, and sparse structures that arise during training (Olsson et al., 2022). These studies provide qualitative and circuit-level insight into model behaviors. Our contribution is to link these geometric structures directly to Bayesian inference in a setting where the posterior is known. We show that keys form near-orthogonal hypothesis axes, queries sharpen onto feasible hypotheses across depth, and value representations unfurl into a one-dimensional entropy manifold. This connects mechanistic interpretability to probabilistic computation in a rigorous way: the internal geometry needed for Bayesian reasoning becomes directly visible.
7.4. Architectural Comparisons and the Copying Problem
[…]
@@ 7.5. Training Dynamics
Finally, concurrent work analyzes the gradient dynamics that create these structures during training (Agarwal et al., 2025). They show that attention and value updates follow coupled laws that produce a stable routing frame and a progressively refined value manifold. Our empirical findings align with this picture: attention stabilizes early, while value vectors continue to encode the posterior with increasing resolution. Together, these perspectives connect the optimization trajectory to the geometric structure that implements Bayesian inference.
@@ 8. Limitations and Future Work
Bijections and HMMs capture essential elements of Bayesian computation–discrete elimination and recursive state tracking–but they represent only a narrow slice of the inference problems encountered by large language models. Future wind tunnels could incorporate richer latent-variable structures, including Kalman filtering, hierarchical Bayesian models, or causal graphical models, all of which have closed-form posteriors and allow precise verification.
@@
Although the hypothesis spaces in both tasks are large enough to prevent memorization, their representational dimensionality is modest (e.g., five hidden states in HMMs). Larger systems with high-dimensional latent variables would test whether the geometric mechanisms we observe–orthogonal hypothesis axes, progressive Q–K sharpening, and value-manifold refinement–scale smoothly with dimensionality.
@@
The experiments here use standard transformers. It remains unclear whether alternative architectures—state-space models, deep MLPs with more sophisticated gating, or hybrid recurrent-attention systems—can form comparable Bayesian manifolds. Wind-tunnel evaluations could provide a principled benchmark for comparing architectures in terms of inference fidelity rather than perplexity alone.
@@
A notable empirical phenomenon is the frame–precision dissociation: attention maps stabilize early while value manifolds continue to unfurl and refine posterior precision. A systematic study of these phases—how early the frame forms, how quickly precision improves, and how these dynamics depend on depth, width, and data complexity—could lead to a more general theory of representation formation in transformers.
@@ 9. Conclusion
We introduced Bayesian wind tunnels—controlled experimental settings with analytic posteriors and combinatorially large hypothesis spaces—to test whether neural sequence models genuinely implement Bayesian inference rather than merely mimicking it. Across multiple inference problems, small transformers converge to the exact Bayesian posterior with sub-bit calibration error, even at sequence lengths well beyond those seen in training. The key insight is that Bayesian inference is not monolithic. We decompose it into three inference primitives—belief accumulation, belief transport, and random-access binding—and show that different architectures realize different subsets:
  • Transformers realize all three primitives and succeed uniformly.
  • Mamba realizes accumulation and transport, achieving state-of-the-art on HMM filtering, but struggles with random-access binding.
  • LSTMs realize only accumulation of static sufficient statistics: they succeed on belief revision (where the statistic is fixed-dimensional) but fail when the statistic must evolve under dynamics or be indexed by content.
  • MLPs realize none and fail uniformly.
Geometric diagnostics reveal how these primitives are implemented. Keys form an approximately orthogonal basis over hypotheses; queries progressively align with the feasible region of that basis; and value vectors organize along a low-dimensional manifold parameterized by posterior entropy. On HMM tracking, Mamba’s representations organize into five discrete clusters—one per hidden state—showing that the model discovers the corner geometry of the belief simplex. Training sculpts this manifold: attention patterns stabilize early, while value representations continue refining posterior precision—a frame–precision dissociation predicted by concurrent gradient-dynamics analysis. The wind-tunnel regime is intentionally simplified, but it establishes a clear lower bound: if a model cannot implement Bayes in settings where the posterior is known and memorization is impossible, it cannot do so in natural language. Conversely, our results show that content-based routing is sufficient for exact Bayesian inference when the task demands only the primitives that architecture can realize. This provides a principled foundation for studying approximate reasoning in larger models and offers concrete, testable predictions—orthogonal hypothesis axes, progressive Q–K sharpening, and value-manifold structure—for analysing pretrained LLMs. The primitives framework explains why transformers succeed: they furnish the architectural mechanisms for all three inference primitives. Attention provides random-access binding through query-key matching; content-based routing enables belief transport; and the residual stream accumulates evidence across positions. Understanding how these primitives scale to real-world language, and whether new architectures can realize them more efficiently, remains an important direction for future work.